5. Are the following linearly independent? Show your work. x; = 3t+2, x, = 2te2+, x= t + 3e2t
Answer #1

3 h = t +3e2t not to zero. © Given, x= 3t+2 = at et { 2, was} in Linearly independent if their wronskiau is equal Now, W (2,X